OpenStudy (anonymous):

PLEASE HELP! Which of the following is an equivalent form of the compound inequality −44 > −2x − 8 >=−8? −2x − 8 > −44 and −2x − 8 >=−8 −2x > −44 and −8 >=−8 −2x − 8 < −44 and −2x − 8 =<−8 −2x − 8 < −44 and −2x − 8 =<−8

OpenStudy (anonymous):

add 8 divide by -2 when you divide change all the inequalities. i.e. flip them so \[\geq\] becomes \[\leq\] and so on

OpenStudy (anonymous):

\[−44 > −2x − 8 \geq−8\] \[-44+8>-2x-8+8\geq-8+8\] \[-36>-2x\geq 0\] is there a typo here?

OpenStudy (anonymous):

because there is no way that this has a solution. the very first line has \[-44>-8\] which is certainly not true. so you cannot solve this

OpenStudy (anonymous):

I've gotten that a lot. I think it's just the math program I am in. sorry :/

OpenStudy (anonymous):

you should tell your teacher that it makes no sense to start out with an inequality that looks like \[-44>\text{expression } >-8\] since it is wrong at the start!

OpenStudy (anonymous):

is this an on line program? love to know which one. really.